# Adaptive Test-Time Training for Predicting Need for Invasive Mechanical Ventilation in Multi-Center Cohorts

### v1

@@ 1 Introduction
Invasive mechanical ventilation (IMV) is a critical intervention utilized in intensive care units (ICUs) for patients with severe respiratory failure and acute respiratory distress syndrome (ARDS) grotberg2023management. However, its use is complicated by the risk of ventilator-induced lung injury and complications resulting from prolonged IMV. Timely and accurate identification of patients at high risk for mechanical ventilation is crucial for optimizing clinical decision-making. Early recognition of these patients enables proactive medical interventions and facilitates efficient resource allocation within hospital system (fan2018acute).
@@
In recent years, the development of machine learning (ML) models has shown great promise in predicting the need for IMV, which leverages electronic health record (EHR) data to identify complex patterns that human clinicians might overlook (shashikumar2021development). These models can incorporate diverse features, including vital signs and laboratory results to enhance prediction accuracy and provide critical decision support in ICU settings. However, the effective deployment of such models in real-world clinical settings remains a challenge. A key issue is the variability in data distributions across hospitals due to differences in patient populations, clinical practices, and EHR systems. These shifts, often referred to as domain shifts, can substantially degrade the performance of predictive models that were trained on data from a single or limited number of sources. For instance, a respiratory failure prediction model (lam2024development) trained on ICU cohort from UC San Diego Health showed an approximately 12% drop in the area under the curve (AUC) when evaluated on an external ICU cohort.
@@
Addressing this challenge requires adaptive methodologies that can account for site-specific heterogeneity. Existing approaches include pre-training models on large multi-center datasets (shashikumar2021development), and transfer learning to fine-tune models on site-specific data (lam2024development) to align feature representations across domains. While these methods have shown promise, many require access to labeled data from the target domain during training or involve computationally expensive retraining processes, which are not always feasible in real-time clinical settings. Test-time training (TTT) offers a novel and efficient solution to this problem by enabling models to adapt dynamically at the time of prediction, without requiring pre-access to target domain labels or costly re-training. TTT leverages an auxiliary task, trained alongside the main task, to update the model’s parameters or representations using the test input itself.
@@
Predictive systems based on EHR data, such as Composer (shashikumar2021artificial), have shown significant real-world impact to improve clinical outcomes through real-time decision support (boussina2024impact). In a before-and-after quasi-experimental design study at two emergency departments (EDs), the Composer model for sepsis prediction significantly increased bundle compliance and reduced in-hospital mortality. However, limited prior work has explored TTT in the context of EHR data, particularly in real-world clinical scenarios. By enabling dynamic adaptation at prediction time, TTT addresses variability across institutions and patient populations, ensuring robust performance in critical tasks, such as predicting IMV need in multi-center cohorts. Unlike pretraining/offline alignment approaches (e.g., MaskTab (chenmasktab), PhyMask (kara2024phymask), and SPOT (gurumoorthy2021spot)), which require source data and extended training, our setting is source-free and demands on-the-fly adaptation at deployment.
@@ 2.1 Predictive Models for Invasive Mechanical Ventilation
Early IMV-risk tools such as ROX and regression scores are interpretable but struggle with nonlinear, time-varying physiology (roca2019index). Leveraging EHR-scale data, VentNet predicts IMV 24h ahead with a feedforward model (shashikumar2021development); encoder–decoder designs like DBNet integrate structured signals and demographics (zhang2021dbnet); and multimodal hybrids that fuse CXR with EHR further boost discrimination (tandon2023hybrid). However, cross-site performance often degrades due to population, workflow, and EHR heterogeneity; recovery via target-domain fine-tuning is common but label-intensive and operationally impractical for continuous deployment.
@@ 2.2 Test-Time Adaptation
TTA adapts models on unlabeled test inputs without revisiting source data (liang2024comprehensive). Batch normalization(BN)-centric methods include prediction-time BN statistics updates (nado2020evaluating) and TENT’s entropy minimization for BN parameters (wang2020tent), while source-free SHOT freezes the classifier and adapts the encoder with pseudo-labels (liang2020we). Test-time training (TTT) attaches auxiliary SSL branches for online encoder updates (sun2020test); extensions like TTT++ (contrastive) and ClusT3 (clustering) improve alignment but may inherit instability or assume domain consistency (liu2021ttt++; hakim2023clust3).
Three relevant directions are: (i) T3A, an optimization-free method forming class prototypes from streaming test data to reweight logits that is efficient but classifier-level only, assuming stable class structure (iwasawa2021test); (ii) SAR, which filters unreliable samples and applies sharpness-aware entropy minimization for stable BN updates that is effective with small batches yet BN-dependent (niu2023towards); (iii) CoTTA, maintaining a moving teacher with augmentation- and weight-averaged pseudo-labels plus periodic restoration, is useful for long horizons but hyperparameter-sensitive with potential error accumulation (wang2022continual).
@@ 3.1 Preliminary: Test-Time Training
Test-Time Training (sun2020test) addresses these shifts by leveraging the same SSL task during both the training and testing phases to align features between the training domain and individual test instances. The framework consists of a shared feature encoder $f_{e}(\cdot;\theta_{e})$, a primary classification head $h_{c}(\cdot;\theta_{c})$ and an SSL head $h_{s}(\cdot;\theta_{s})$.
@@ 3.2 Theoretical Insights
Prior work (liu2021ttt++) derives accuracy bounds under assumptions of distributional alignment and task consistency. We provide an independent perspective based on information theory that examines how the auxiliary task informs the main task through shared representations. Let $Z$ and $Z^{\prime}$ represent the feature representations from the feature encoder for the training and test domains, respectively. We define $\pi_{m}$ is the main task classifier predicting $Y_{m}$, and $\pi_{s}$ is the SSL classifier predicting $Y_{s}$. The probability $P(\pi_{m}(Z^{\prime})=Y_{m}^{\prime})$ quantifies the likelihood that the main task classifier $\pi_{m}$ correctly predicts the main task label $Y_{m}^{\prime}$ at test time.
@@ 3.3.2 Prototype-Guided Adaptation
Test-Time Training Stage. Traditional Optimal Transport (OT) assumes a full alignment between source and target distributions, which may be too rigid in the presence of domain shifts. We refine the alignment between the training prototypes $\mathbf{P}$ and the test-time feature representations $z^{\prime}$ by incorporating POT. Instead of constraining the transport plan to only partially align prototypes with the test instance (chapel2020partial), we augment the set of $\mathbf{z}^{\prime}$ by adding k1k-1 perturbed duplicates to transform the transport problem into a standard optimal transport setting while enabling partial matches between $z^{\prime}$ and the prototypes $\mathbf{P}$.
@@ 4.1 Experimental Setting
Implementation Details. Our network architecture follows Vent.io (lam2024development) (Appendix B.2). We use Bayesian optimization for source-domain pretraining to tune general network hyperparameters (learning rate, weight regularization, number of hidden layers). The prototype set size is $k=4$. For dynamic masking, we adopt a two-phase schedule: a warm-up phase with a fixed prior mask probability of $0.5$, followed by an adaptive phase where masking probabilities are updated from feature relevance. During deployment, test-time training (TTT) performs five gradient steps per input (we refer to each step as an “iteration”), and we follow the standard reset protocol (sun2020test) by restoring encoder weights to the pretrained state after each instance. For partial optimal transport, we use Sinkhorn with entropic regularization $\varepsilon=0.1$, a maximum of 1000 Sinkhorn iterations, and a mini-batch size equal to the prototype count ($K=k$). All baselines are re-implemented with the same encoder, data pipeline, and hyperparameter search budget to ensure a fair comparison.
Evaluation. We follow the clinical labeling scheme (lam2024development) to capture the physiological states of respiratory failure, defining score $\geq\!3$ as positive and <3<\!3 as control (Appendix B.1). Encounters are categorized as True Positive, False Positive, True Negative, or False Negative based on predictions within the specified prediction window (criteria in Appendix B.1). Performance is reported as AUC with mean $\pm$ standard error over 20 independent test-time training runs. In addition, following best practices for clinical prediction model assessment (huang2020tutorial), we report Brier score for all models to evaluate calibration and clinical utility more comprehensively.
@@ 4.2 Comparison with Baselines
Baselines. We consider a set of foundational and representative TTT and TTA methods and adapt each to the EHR domain to ensure fair and meaningful comparison in our experimental evaluations (Appendix B.3 provides full details of adapting baselines to EHR setting): TEST, TENT (wang2020tent), TTT (sun2020test), TTT++ (liu2021ttt++), ClusT3 (hakim2023clust3), NC-TTT (osowiechi2024nc), T3A (iwasawa2021test), SAR (niu2023towards) and CoTTA (wang2022continual). We also evaluate two ablated versions of our method: PriTTT, which removes the adaptive distribution matching module and relies solely on updates to the mask probabilities for test-time training. DynTTT, which removes the dynamic masking module and focuses only on adaptive distribution matching.
@@
Computational Cost. Our proposed AdaTTT framework incorporates both feature importance updates and optimal transport computation during the test-time training phase. These additional operations inevitably increase the computational cost compared to standard TTT frameworks. We use the Sinkhorn algorithm (sinkhorn1967diagonal) that leverages entropy regularization to ensure scalable computations. We evaluate the execution time of a single gradient update during test-time training. The average execution time is 0.29s, and did not change much, remaining at 0.26s when increasing the prototype size from 4 to 16.
@@ A.2.1 Binary case
Under Hellman’s inequality [hellman1970probability], we have
@@ A.2.2 Multi-class case
Based on the simplified Fano’s inequality [thomas2006elements], we have
@@ B.1 Dataset
Clinical labeling scheme for the various physiological states of respiratory failure. Table 3 lists the labeling criteria used in [lam2024development].
@@ B.3 Baselines
We compare against representative TTA/TTT methods and adapt each fairly to the EHR setting: TEST, TENT [wang2020tent], TTT [sun2020test], TTT++ liu2021ttt++, ClusT3 [hakim2023clust3], NC-TTT [osowiechi2024nc], T3A [iwasawa2021test], SAR [niu2023towards], and CoTTA [wang2022continual].
@@ Methods.
  • TENT [wang2020tent]: minimize prediction entropy at test time; we update only BN affine parameters and running statistics.
  • TTT [sun2020test]: jointly train the network on the main task and the same auxiliary SSL objective as ours (reconstruction + masked–feature modeling) in the source domain; at test time, adapt the encoder only by minimizing this SSL loss. For fairness, TTT uses uniform, feature-agnostic masking (no dynamic masking), and does not use prototypes or optimal transport.
  • TTT++ [liu2021ttt++]: identical SSL setup as TTT above and the same encoder-only test-time updates; additionally aligns first/second-order moments between source and target. Because source data are unavailable at deployment, source moments are cached from the development split during training and used for alignment at test time.
  • ClusT3 [hakim2023clust3]: add a projector on top of the shared encoder and adapt by maximizing mutual information with discrete codes. We use an MLP projector (tabular analogue of the original CNN projector) and the same codebook size across sites.
  • NC-TTT [osowiechi2024nc]: optimize a noise-contrastive auxiliary likelihood at test time. The noise distribution is factorized Gaussian with per-feature mean/variance estimated on the development split.
  • T3A [iwasawa2021test]: optimization-free classifier adjustment that builds class prototypes from confident test predictions and reweights logits. We compute prototypes from the FIFO buffer, with the confidence threshold tuned on the development split; no backprop or encoder change.
  • SAR [niu2023towards]: sharpness-aware entropy minimization with unreliable-sample filtering for small/test-time batches. We follow the BN-only update rule as in TENT, add SAM-style perturbations to BN parameters, and use the same FIFO buffer; filter thresholds are validated once on the development split.
  • CoTTA [wang2022continual]: maintain an EMA teacher and perform augmentation/weight-averaged pseudo-labeling with periodic weight restoration. Tabular augmentation is feature masking; teacher momentum and restoration period are tuned under the shared budget. Encoder is adapted.

### v2

@@ 1 Introduction
Invasive mechanical ventilation (IMV) is a critical intervention utilized in intensive care units (ICUs) for patients with severe respiratory failure and acute respiratory distress syndrome (ARDS) Grotberg et al. (2023). However, its use is complicated by the risk of ventilator-induced lung injury and complications resulting from prolonged IMV. Timely and accurate identification of patients at high risk for mechanical ventilation is crucial for optimizing clinical decision-making. Early recognition of these patients enables proactive medical interventions and facilitates efficient resource allocation within hospital system (Fan et al., 2018).
@@
In recent years, the development of machine learning (ML) models has shown great promise in predicting the need for IMV, which leverages electronic health record (EHR) data to identify complex patterns that human clinicians might overlook (Shashikumar et al., 2021b). These models can incorporate diverse features, including vital signs and laboratory results to enhance prediction accuracy and provide critical decision support in ICU settings. However, the effective deployment of such models in real-world clinical settings remains a challenge. A key issue is the variability in data distributions across hospitals due to differences in patient populations, clinical practices, and EHR systems. These shifts, often referred to as domain shifts, can substantially degrade the performance of predictive models that were trained on data from a single or limited number of sources. For instance, a respiratory failure prediction model (Lam et al., 2024) trained on ICU cohort from UC San Diego Health showed an approximately 12% drop in the area under the curve (AUC) when evaluated on an external ICU cohort.
@@
Addressing this challenge requires adaptive methodologies that can account for site-specific heterogeneity. Existing approaches include pre-training models on large multi-center datasets (Shashikumar et al., 2021b), and transfer learning to fine-tune models on site-specific data (Lam et al., 2024) to align feature representations across domains. While these methods have shown promise, many require access to labeled data from the target domain during training or involve computationally expensive retraining processes, which are not always feasible in real-time clinical settings. Test-time training (TTT) offers a novel and efficient solution to this problem by enabling models to adapt dynamically at the time of prediction, without requiring pre-access to target domain labels or costly re-training. TTT leverages an auxiliary task, trained alongside the main task, to update the model’s parameters or representations using the test input itself.
@@
Predictive systems based on EHR data, such as Composer (Shashikumar et al., 2021a), have shown significant real-world impact to improve clinical outcomes through real-time decision support (Boussina et al., 2024). In a before-and-after quasi-experimental design study at two emergency departments (EDs), the Composer model for sepsis prediction significantly increased bundle compliance and reduced in-hospital mortality. However, limited prior work has explored TTT in the context of EHR data, particularly in real-world clinical scenarios. By enabling dynamic adaptation at prediction time, TTT addresses variability across institutions and patient populations, ensuring robust performance in critical tasks, such as predicting IMV need in multi-center cohorts. Unlike pretraining/offline alignment approaches (e.g., MaskTab (Chen et al., ), PhyMask (Kara et al., 2024), and SPOT (Gurumoorthy et al., 2021)), which require source data and extended training, our setting is source-free and demands on-the-fly adaptation at deployment.
@@ 2.1 Predictive Models for Invasive Mechanical Ventilation
Early IMV-risk tools such as ROX and regression scores are interpretable but struggle with nonlinear, time-varying physiology (Roca et al., 2019). Leveraging EHR-scale data, VentNet predicts IMV 24h ahead with a feedforward model (Shashikumar et al., 2021b); encoder–decoder designs like DBNet integrate structured signals and demographics (Zhang et al., 2021); and multimodal hybrids that fuse CXR with EHR further boost discrimination (Tandon et al., 2023). However, cross-site performance often degrades due to population, workflow, and EHR heterogeneity; recovery via target-domain fine-tuning is common but label-intensive and operationally impractical for continuous deployment.
@@ 2.2 Test-Time Adaptation
TTA adapts models on unlabeled test inputs without revisiting source data (Liang et al., 2024). Batch normalization(BN)-centric methods include prediction-time BN statistics updates (Nado et al., 2020) and TENT’s entropy minimization for BN parameters (Wang et al., 2020), while source-free SHOT freezes the classifier and adapts the encoder with pseudo-labels (Liang et al., 2020). Test-time training (TTT) attaches auxiliary SSL branches for online encoder updates (Sun et al., 2020); extensions like TTT++ (contrastive) and ClusT3 (clustering) improve alignment but may inherit instability or assume domain consistency (Liu et al., 2021; Hakim et al., 2023).
Three relevant directions are: (i) T3A, an optimization-free method forming class prototypes from streaming test data to reweight logits that is efficient but classifier-level only, assuming stable class structure (Iwasawa and Matsuo, 2021); (ii) SAR, which filters unreliable samples and applies sharpness-aware entropy minimization for stable BN updates that is effective with small batches yet BN-dependent (Niu et al., 2023); (iii) CoTTA, maintaining a moving teacher with augmentation- and weight-averaged pseudo-labels plus periodic restoration, is useful for long horizons but hyperparameter-sensitive with potential error accumulation (Wang et al., 2022).
@@ 3.1 Preliminary: Test-Time Training
Test-Time Training (Sun et al., 2020) addresses these shifts by leveraging the same SSL task during both the training and testing phases to align features between the training domain and individual test instances. The framework consists of a shared feature encoder $f_{e}(\cdot;\theta_{e})$, a primary classification head $h_{c}(\cdot;\theta_{c})$ and an SSL head $h_{s}(\cdot;\theta_{s})$.
@@ 3.2 Theoretical Insights
Prior work (Liu et al., 2021) derives accuracy bounds under assumptions of distributional alignment and task consistency. We provide an independent perspective based on information theory that examines how the auxiliary task informs the main task through shared representations. Let $Z$ and $Z^{\prime}$ represent the feature representations from the feature encoder for the training and test domains, respectively. We define $\pi_{m}$ is the main task classifier predicting $Y_{m}$, and $\pi_{s}$ is the SSL classifier predicting $Y_{s}$. The probability $P(\pi_{m}(Z^{\prime})=Y_{m}^{\prime})$ quantifies the likelihood that the main task classifier $\pi_{m}$ correctly predicts the main task label $Y_{m}^{\prime}$ at test time.
@@ 3.3.2 Prototype-Guided Adaptation
Test-Time Training Stage. Traditional Optimal Transport (OT) assumes a full alignment between source and target distributions, which may be too rigid in the presence of domain shifts. We refine the alignment between the training prototypes $\mathbf{P}$ and the test-time feature representations $z^{\prime}$ by incorporating POT. Instead of constraining the transport plan to only partially align prototypes with the test instance (Chapel et al., 2020), we augment the set of $\mathbf{z}^{\prime}$ by adding k1k-1 perturbed duplicates to transform the transport problem into a standard optimal transport setting while enabling partial matches between $z^{\prime}$ and the prototypes $\mathbf{P}$.
@@ 4.1 Experimental Setting
Implementation Details. Our network architecture follows Vent.io (Lam et al., 2024) (Appendix B.2). We use Bayesian optimization for source-domain pretraining to tune general network hyperparameters (learning rate, weight regularization, number of hidden layers). The prototype set size is $k=4$. For dynamic masking, we adopt a two-phase schedule: a warm-up phase with a fixed prior mask probability of $0.5$, followed by an adaptive phase where masking probabilities are updated from feature relevance. During deployment, test-time training (TTT) performs five gradient steps per input (we refer to each step as an “iteration”), and we follow the standard reset protocol (Sun et al., 2020) by restoring encoder weights to the pretrained state after each instance. For partial optimal transport, we use Sinkhorn with entropic regularization $\varepsilon=0.1$, a maximum of 1000 Sinkhorn iterations, and a mini-batch size equal to the prototype count ($K=k$). All baselines are re-implemented with the same encoder, data pipeline, and hyperparameter search budget to ensure a fair comparison.
Evaluation. We follow the clinical labeling scheme (Lam et al., 2024) to capture the physiological states of respiratory failure, defining score $\geq\!3$ as positive and <3<\!3 as control (Appendix B.1). Encounters are categorized as True Positive, False Positive, True Negative, or False Negative based on predictions within the specified prediction window (criteria in Appendix B.1). Performance is reported as AUC with mean $\pm$ standard error over 20 independent test-time training runs. In addition, following best practices for clinical prediction model assessment (Huang et al., 2020), we report Brier score for all models to evaluate calibration and clinical utility more comprehensively.
@@ 4.2 Comparison with Baselines
Baselines. We consider a set of foundational and representative TTT and TTA methods and adapt each to the EHR domain to ensure fair and meaningful comparison in our experimental evaluations (Appendix B.3 provides full details of adapting baselines to EHR setting): TEST, TENT (Wang et al., 2020), TTT (Sun et al., 2020), TTT++ (Liu et al., 2021), ClusT3 (Hakim et al., 2023), NC-TTT (Osowiechi et al., 2024), T3A (Iwasawa and Matsuo, 2021), SAR (Niu et al., 2023) and CoTTA (Wang et al., 2022). We also evaluate two ablated versions of our method: PriTTT, which removes the adaptive distribution matching module and relies solely on updates to the mask probabilities for test-time training. DynTTT, which removes the dynamic masking module and focuses only on adaptive distribution matching.
@@
Computational Cost. Our proposed AdaTTT framework incorporates both feature importance updates and optimal transport computation during the test-time training phase. These additional operations inevitably increase the computational cost compared to standard TTT frameworks. We use the Sinkhorn algorithm (Sinkhorn, 1967) that leverages entropy regularization to ensure scalable computations. We evaluate the execution time of a single gradient update during test-time training. The average execution time is 0.29s, and did not change much, remaining at 0.26s when increasing the prototype size from 4 to 16.
@@ A.2.1 Binary case
Under Hellman’s inequality [Hellman and Raviv, 1970], we have
@@ A.2.2 Multi-class case
Based on the simplified Fano’s inequality [Thomas and Joy, 2006], we have
@@ B.1 Dataset
Clinical labeling scheme for the various physiological states of respiratory failure. Table 3 lists the labeling criteria used in [Lam et al., 2024].
@@ B.3 Baselines
We compare against representative TTA/TTT methods and adapt each fairly to the EHR setting: TEST, TENT [Wang et al., 2020], TTT [Sun et al., 2020], TTT++ Liu et al. [2021], ClusT3 [Hakim et al., 2023], NC-TTT [Osowiechi et al., 2024], T3A [Iwasawa and Matsuo, 2021], SAR [Niu et al., 2023], and CoTTA [Wang et al., 2022].
@@ Methods.
  • TENT [Wang et al., 2020]: minimize prediction entropy at test time; we update only BN affine parameters and running statistics.
  • TTT [Sun et al., 2020]: jointly train the network on the main task and the same auxiliary SSL objective as ours (reconstruction + masked–feature modeling) in the source domain; at test time, adapt the encoder only by minimizing this SSL loss. For fairness, TTT uses uniform, feature-agnostic masking (no dynamic masking), and does not use prototypes or optimal transport.
  • TTT++ [Liu et al., 2021]: identical SSL setup as TTT above and the same encoder-only test-time updates; additionally aligns first/second-order moments between source and target. Because source data are unavailable at deployment, source moments are cached from the development split during training and used for alignment at test time.
  • ClusT3 [Hakim et al., 2023]: add a projector on top of the shared encoder and adapt by maximizing mutual information with discrete codes. We use an MLP projector (tabular analogue of the original CNN projector) and the same codebook size across sites.
  • NC-TTT [Osowiechi et al., 2024]: optimize a noise-contrastive auxiliary likelihood at test time. The noise distribution is factorized Gaussian with per-feature mean/variance estimated on the development split.
  • T3A [Iwasawa and Matsuo, 2021]: optimization-free classifier adjustment that builds class prototypes from confident test predictions and reweights logits. We compute prototypes from the FIFO buffer, with the confidence threshold tuned on the development split; no backprop or encoder change.
  • SAR [Niu et al., 2023]: sharpness-aware entropy minimization with unreliable-sample filtering for small/test-time batches. We follow the BN-only update rule as in TENT, add SAM-style perturbations to BN parameters, and use the same FIFO buffer; filter thresholds are validated once on the development split.
  • CoTTA [Wang et al., 2022]: maintain an EMA teacher and perform augmentation/weight-averaged pseudo-labeling with periodic weight restoration. Tabular augmentation is feature masking; teacher momentum and restoration period are tuned under the shared budget. Encoder is adapted.